OpenStudy (anonymous):

A square with sides of x + 2 and an equilateral triangle of sides with length 2x have equal perimeters. x = ?

OpenStudy (anonymous):

equilateral triangle means that each of the triangles three sides is equal, the squares sides are also all equal as pre definition of a square. a triangle has three sides a square has four sides sum of squares sides = 4(x+2) sum of triangles sides = 3(2x) their perimiters are equal 4(x+2)=3(2x) just algebra from there